OpenStudy (anonymous):

Three consecutive odd integers have a sum of 45. Find the integers

OpenStudy (anonymous):

@freckles

OpenStudy (vera_ewing):

So you know that the integers have to be odd, right?

OpenStudy (anonymous):

Ya ..

OpenStudy (vera_ewing):

And they have to be consecutive. (1, 2, 3)

OpenStudy (vera_ewing):

So just try it! Start with 13.

OpenStudy (vera_ewing):

What is 13+15+17?

OpenStudy (anonymous):

Lol 😐😐😐😑😑 I'm dumb

OpenStudy (vera_ewing):

It's simple addition. Use a calculator if you absolutely have to.

OpenStudy (anonymous):

Can you help me with some more in a bit if I message you?

OpenStudy (vera_ewing):

I'll try.

OpenStudy (anonymous):

Thanks babe!

OpenStudy (jhannybean):

consecutive odd integers: have a difference of 2 between each of the numbers, and theyre odd, aka 1,3,5... etc \[2n + (2n+1) +(2n+3)=45\]

OpenStudy (freckles):

I would solve the equation n+(n+2)+(n+4)=45 but I think @Jhannybean meant to say (2n-1)+(2n+1)+(2n+3)=45 which is also a great route to go :)

OpenStudy (jhannybean):

Oh that's right, thanks @freckles :D